LIFETIME-MEDIATED POLARIZATION EFFECTS IN NONLINEAR SPECTROSCOPY: DEGENERATE FOUR-WAVE MIXING STUDIES OF PREDISSOCIATED S_{2}O IN A SLIT-JET EXPANSION

\underline{THOMAS MÜLLER}, PATRICK DUPRÉ^{a}, QINGGUO ZHANG^{b}, and PATRICK H. VACCARO, Department of Chemistry, Yale University, 225 Prospect Street, New Haven, CT 06520.

Transient S_{2}O molecules were entrained in a pulsed slit-jet expansion (T_{rot} < 10 K) and interrogated through use of sub-Doppler \underline{D}egenerate \underline{F}our-\underline{W}ave \underline{M}ixing (DFWM) spectroscopy. High-resolution scans have been acquired for the 2_{o}^{v} (v=3-10) vibronic bands of the intense C̃ ^{1}A′ ← X̃ ^{1}A′ (π^{∗} ←π) absorption system, where increasing excitation of the ν_{2} S-S stretching mode is known to promote predissociation of the C̃ state. For bands involving moderately-predissociated states (e.g., 2_{0}^{4} where τc̃ ≃ 63 ps), the recorded pattern of rovibronic line intensities exhibits a pronounced dependence upon DFWM polarization geometry, a situation not encountered in analogous studies performed for features terminating on long-lived levels of the C̃ manifold (e.g., 2_{0}^{3} where τc̃ ≃ 22 ns). This behavior can be reproduced quantitatively through detailed weak-field analyses of the resonant DFWM response, however, a qualitative understanding follows from the selective dissipation of optically-induced transient gratings as incurred by unimolecular relaxation pathways. In strongly-predissociated members of the 2_{0}^{v} progression ( i.e., v ≥ 5 ), additional polarization specificity is introduced by the presence of strong depopulation pumping processes which lead to the creation of net orientation and/or alignment of the molecular ensemble on a timescale commensurate with that of the pulsed four-wave mixing experiment Owing to its absorption-based response and laser-limited spectral resolution, DFWM is often applied to target species where rapid nonraditive relaxation channels preclude successful exploitation of detection techniques based upon secondary matter-field interactions (e.g., fluorescence or ionization). Therefore, a detailed understanding of the role which molecular lifetime plays as a mediator for resonant nonlinear response is of central importance for the quantitative application of this optical scheme.

aPermanent address: High Magnetic Field Laboratory, CNRS, BP 166, 38042 Grenoble, Cedex 9 (France)

bPresent address: G. R. Harrison Spectroscopy Laboratory, MIT, Cambridge, MA 02139